OpenStudy (anonymous):

8(9^x)+3(6^x)-81(4^x)=0 help pleaaase :D

OpenStudy (anonymous):

divide by 6^x

OpenStudy (klimenkov):

Or by \(4^x\)

OpenStudy (anonymous):

\[8\times(\frac{9}{6})^x+3-81\times(\frac{4}{6})^x=0\]\[8\times(\frac{3}{2})^x+3-81\times(\frac{2}{3})^x=0\]and let \[t=(\frac{3}{2})^x\]

OpenStudy (anonymous):

make a quadratic in terms of t and solve for t

OpenStudy (anonymous):

and also u will have a similar process with what klimenkov said